# arXiv:2603.01669v1[math.CO] 02 Mar 2026

### v1

@@ 1 Introduction
An overpartition of $n$ extends the unrestricted partitions by allowing the first occurrence of each part size to be overlined. The number of overpartitions of $n$ is denoted by $\bar{p}(n)$. For instance, the overpartition of $3$ are
@@
Recently, Paksok and Saikia [10] proved new Ramanujan-type congruence modulo $8$ and $16$ for $\bar{a}_{2m+1}^{*}(n)$, $\bar{a}_{2m+2}^{*}(n)$, and $\bar{a}_{8m+3}^{*}(n)$. Later, Das et. al [6] have studied the divisibility properties of $\bar{a}_{r}^{*}(n)$ for different values of $r$. In particular, they proved the following generalizations of Seller’s results [11, Th. 2.5 and Th. 2.7].
@@
It is interesting to note, the generating function (1.4) unifies (1.2) and (1.3). In this paper, we prove several infinite families of congruences modulo prime and modulo powers of $2$ for $\bar{a}_{r,s}(n)$ which generalizes Theorem 1.1 and 1.2. The main results of this paper pertains to several Ramanujan-type congruences modulo powers of $2$ and modulo prime $p\geq 3$, satisfied by $\bar{a}_{r,s}$.
@@
In the next two theorems we prove infinite families of congruence modulo powers of $2$.
@@ 4 Congruences modulo powers of 22
We first note that for $s=2i$ in Theorem 1.3, if $3n+1=k^{2}$, then we have the divisibility by $4$. Also, if $3n+1=2k^{2}$, we have $2\equiv(2k)^{2}\pmod{3}$, which is a contradiction, since $2$ is a quadratic non-residue modulo $3$. This settles (1.8).
Again, for $3n+2=k^{2}$, we have $2\equiv k^{2}\pmod{3}$, which is not true, since $2$ is a quadratic non-residue modulo $3$. Now, it is easy to observe for $r+s=2\ell$, in Theorem 1.3, if $3n+2=2k^{2}$, then we have the divisibility by $4$. This settles (1.9).
Now, for $9n+3=k^{2}$, we note $9n+3=3(3n+1)$ is not divisible by $9$. Thus, $9n+3$ is not a square. Similarly, for $9n+3=2k^{2}$, we have $3(6n+2)=(2k)^{2}$. Since $3(6n+2)$ is not divisible by $9$, implies $18n+6$ is not a square. Therefore, (1.10) follows from Theorem 1.3.
Lastly, for $9n+6=k^{2}$, we observe $9n+6=3(3n+4)$ is not divisible by $9$. Thus, $9n+6$ is not a square. Similarly, for $9n+6=2k^{2}$, we have $3(6n+4)=(2k)^{2}$. Since $3(6n+2)$ is not divisible by $9$, implies $18n+12$ is not square. Therefore, (1.11) follows from Theorem 1.3.
@@
It is easy to observe $9n+3=k^{2}$ and $9n+3=4k^{2}=(2k)^{2}$ are not true, since $9n+3$ is not a square. Similarly, for $9n+3=2(2k)^{2}$ and $9n+3=2(2k-1)^{2}$, we have $3(6n+2)=(8k)^{2}$ and $3(6n+2)=(2(2k-1))^{2}$, respectively which is again not true. Therefore, for any integer $s$ and odd $r$ in Theorem 1.4, if $9n+3=k^{2}+2l^{2}$, then we have the divisibility by $8$. This settles (1.12).
It is easy to see, $9n+6=k^{2}$ and $9n+6=4k^{2}$ are not true. Similarly, for $9n+6=2(2k)^{2}$ and $9n+6=2(2k-1)^{2}$, we have $3(6n+4)=(8k)^{2}$ and $3(6n+4)=(2(2k-1)^{2}$, respectively, which is again not true. Therefore, for any integer $s$ and odd $r$ in Theorem 1.4, if $9n+6=k^{2}+2l^{2}$, then we have the divisibility by $8$. This settles (1.13). ∎
@@ 5 Congruences modulo prime $p\geq 3$
We observe that $3n+2\equiv k^{2}\pmod{3}$ holds if and only if $2\equiv k^{2}\pmod{3}$. Since $2$ is a quadratic non-residue modulo $3$, we obtain
@@
$=\dfrac{1}{f_{p}}\phi(-q^{2})$
$=\dfrac{1}{f_{p}}\sum_{k=-\infty}^{\infty}(-q^{2})^{k^{2}}.$
@@ 6 Concluding remark
$\bar{a}_{2^{k+1}j+2^{k}-1,2i+1}(3n+2)$ $\equiv 0\pmod{2^{k+1}}$ (6.1)
$\bar{a}_{2^{k+1}j+2^{k}-1,2i+1}(9n+3)$ $\equiv 0\pmod{2^{k+2}}$ (6.2)
$\bar{a}_{2^{k+1}j+2^{k}-1,2i+1}(9n+6)$ $\equiv 0\pmod{2^{k+2}}$. (6.3)

### v2

# Overcolored Partition Restricted by Parity of the Parts

@@ 1 Introduction
An overpartition of $n$ extends the unrestricted partitions by allowing the first occurrence of each part size to be overlined. The number of overpartitions of $n$ is denoted by $\bar{p}(n)$. For instance, the overpartition of 33 are
@@
Recently, Paksok and Saikia [10] proved new Ramanujan-type congruence modulo 88 and 1616 for $\bar{a}_{2m+1}^{*}(n)$, $\bar{a}_{2m+2}^{*}(n)$, and $\bar{a}_{8m+3}^{*}(n)$. Later, Das et. al [6] have studied the divisibility properties of $\bar{a}_{r}^{*}(n)$ for different values of $r$. In particular, they proved the following generalizations of Seller’s results [11, Th. 2.5 and Th. 2.7].
@@
It is interesting to note, the generating function (1.4) unifies (1.2) and (1.3). In this paper, we prove several infinite families of congruences modulo prime and modulo powers of 22 for $\bar{a}_{r,s}(n)$ which generalizes Theorem 1.1 and 1.2. The main results of this paper pertains to several Ramanujan-type congruences modulo powers of 22 and modulo prime $p\geq 3$, satisfied by $\bar{a}_{r,s}$.
@@
In the next two theorems we prove infinite families of congruence modulo powers of 22.
@@ 4 Congruences modulo powers of 22
We first note that for $s=2i$ in Theorem 1.3, if $3n+1=k^{2}$, then we have the divisibility by 44. Also, if $3n+1=2k^{2}$, we have $2\equiv(2k)^{2}\pmod{3}$, which is a contradiction, since 22 is a quadratic non-residue modulo 33. This settles (1.8).
Again, for $3n+2=k^{2}$, we have $2\equiv k^{2}\pmod{3}$, which is not true, since 22 is a quadratic non-residue modulo 33. Now, it is easy to observe for $r+s=2\ell$, in Theorem 1.3, if $3n+2=2k^{2}$, then we have the divisibility by 44. This settles (1.9).
Now, for $9n+3=k^{2}$, we note $9n+3=3(3n+1)$ is not divisible by 99. Thus, $9n+3$ is not a square. Similarly, for $9n+3=2k^{2}$, we have $3(6n+2)=(2k)^{2}$. Since $3(6n+2)$ is not divisible by 99, implies $18n+6$ is not a square. Therefore, (1.10) follows from Theorem 1.3.
Lastly, for $9n+6=k^{2}$, we observe $9n+6=3(3n+4)$ is not divisible by 99. Thus, $9n+6$ is not a square. Similarly, for $9n+6=2k^{2}$, we have $3(6n+4)=(2k)^{2}$. Since $3(6n+2)$ is not divisible by 99, implies $18n+12$ is not square. Therefore, (1.11) follows from Theorem 1.3.
@@
It is easy to observe $9n+3=k^{2}$ and $9n+3=4k^{2}=(2k)^{2}$ are not true, since $9n+3$ is not a square. Similarly, for $9n+3=2(2k)^{2}$ and $9n+3=2(2k-1)^{2}$, we have $3(6n+2)=(8k)^{2}$ and $3(6n+2)=(2(2k-1))^{2}$, respectively which is again not true. Therefore, for any integer $s$ and odd $r$ in Theorem 1.4, if $9n+3=k^{2}+2l^{2}$, then we have the divisibility by 88. This settles (1.12).
It is easy to see, $9n+6=k^{2}$ and $9n+6=4k^{2}$ are not true. Similarly, for $9n+6=2(2k)^{2}$ and $9n+6=2(2k-1)^{2}$, we have $3(6n+4)=(8k)^{2}$ and $3(6n+4)=(2(2k-1)^{2}$, respectively, which is again not true. Therefore, for any integer $s$ and odd $r$ in Theorem 1.4, if $9n+6=k^{2}+2l^{2}$, then we have the divisibility by 88. This settles (1.13). ∎
@@ 5 Congruences modulo prime $p\geq 3$
We observe that $3n+2\equiv k^{2}\pmod{3}$ holds if and only if $2\equiv k^{2}\pmod{3}$. Since 22 is a quadratic non-residue modulo 33, we obtain
@@
$=\dfrac{f_{2p}}{f_{p}^{2}}\phi(-q^{2})$
$=\dfrac{f_{2p}}{f_{p}^{2}}\sum_{k=-\infty}^{\infty}(-q^{2})^{k^{2}}.$
@@ 6 Concluding remark
$\bar{a}_{2^{k+1}j+2^{k}-1,2^{k}i+1}(3n+2)$ $\equiv 0\pmod{2^{k+1}}$ (6.1)
$\bar{a}_{2^{k+1}j+2^{k}-1,2^{k}i+1}(9n+3)$ $\equiv 0\pmod{2^{k+2}}$ (6.2)
$\bar{a}_{2^{k+1}j+2^{k}-1,2^{k}i+1}(9n+6)$ $\equiv 0\pmod{2^{k+2}}$. (6.3)